# MagicQuill: An Intelligent Interactive Image Editing System

### v1

@@ B.1 Failure Case of Editing Processor
Figure 10: Illustration of the Scribble-Prompt Trade-Off. Given user-provided brush strokes (a) with the text prompt “man”, we show generation results with different edge control strengths: (b) with strength of $0.6$ and (c) with strength of $0.2$.
@@
Colorization-Details Trade-Off. We observe a trade-off between colorization accuracy and detail preservation. Since our conditional image inpainting pipeline relies on downsampled color blocks and CNN-extracted edge maps as input, structrual details in the edited regions may be compromised during the generation process.
@@
B.2 Failure Case of Painting Assistor
Ambiguity of the Brush Strokes. Our system enables users to express their editing intentions through brush strokes, which are then interpreted by the Painting Assistor via Draw&Guess. However, this approach faces inherent limitations due to the ambiguous nature of user-provided sketches. For instance, a simple circular sketch could represent various objects like strawberry, raspberry, or candy, making it challenging for the model to accurately infer the user’s intended modification, as shown in Fig. 12.
@@ B.2 Failure Case of Painting Assistor
Appendix C Generalizability of Editing Processor
@@
Appendix D In-Context Editing Intent Interpretation
@@ Appendix E User Study Details and Questionnaires
To control for learning effects, we randomly divided participants into two groups: Group A used MagicQuill before the baseline (Fig. 15), while Group B followed the reverse order. Each participant completed a comprehensive evaluation consisting of $10$ questions per system, modified from the System Usability Scale (SUS) [5], spanning four key categories: Complexity and Efficiency, Consistency and Integration, Ease of Use, and Overall Satisfaction. . The detailed evaluation results are presented in Fig. 16. Additionally, participants responded to $2$ specific questions addressing the Painting Assistor’s accuracy and efficiency detailed in Sec. 4.2.

### v2

@@ 10.1 Failure Case of Editing Processor
Figure 15: Illustration of the Scribble-Prompt Trade-Off. Given user-provided brush strokes (a) with the text prompt “man”, we show generation results with different edge control strengths: (b) with edge control of $0.6$ and (c) with edge control of $0.2$.
@@
Colorization-Details Trade-Off: We observe a trade-off between colorization accuracy and detail preservation. Since our conditional image inpainting pipeline relies on downsampled color blocks and CNN-extracted edge maps as input, structrual details in the edited regions may be compromised during the generation process.
@@
10.2 Failure Case of Painting Assistor
Ambiguity of the Sketch: Our system enables users to express their editing intentions through brush strokes, which are then interpreted by the Painting Assistor via Draw&Guess. However, this approach faces inherent limitations due to the ambiguous nature of user-provided sketches. For instance, a simple circular sketch could represent various objects like strawberry, raspberry, or candy, making it challenging for the model to accurately infer the user’s intended modification, as shown in Fig. 17.
@@ 10.2 Failure Case of Painting Assistor
11 Generalizability of Editing Processor
@@
12 In-Context Editing Intent Interpretation
@@ 13 User Study Details and Questionnaires
To control for learning effects, we randomly divided participants into two groups: Group A used MagicQuill before the baseline (Fig. 21), while Group B followed the reverse order. Each participant completed a comprehensive evaluation consisting of $10$ questions per system, modified from the System Usability Scale (SUS) [7], spanning four key categories: Complexity and Efficiency, Consistency and Integration, Ease of Use, and Overall Satisfaction. . The detailed evaluation results are presented in Fig. 20. Additionally, participants responded to $2$ specific questions addressing the Painting Assistor’s accuracy and efficiency.